Consider a container with 9.05g of N2 gas and an unknown mass of CO2 gas. If we know that the partial pressure of the CO2 is twice that of the N2, determine the mass of CO2 gas in the container.
1
Expert's answer
2020-09-17T07:20:42-0400

P is proportional to n

Pco2/PN2 = nco2/nN2

Moles of gas is equal to the weight/molecular weight of gas

Pco2 is twice that of PN2

Pco2 = 2PN2, weight of N2 = 9.05 gram

2PN2/PN2 = nco2/nN2

2 nN2 = nco2

2w/m.w (N2) = w/m.w (co2)

2×9.05/14 = w co2 /44

Wco2 = (2×9.05×44)/14

Wco2 = 56.88 gram